OpenStudy (anonymous):

Find the domain of the function. g(x)=^4sq.rt x^2+2x

OpenStudy (anonymous):

yep that's what it's supposed to be

OpenStudy (irishboy123):

scan or link the question pls

OpenStudy (irishboy123):

sorry, been here done that! just scan or link it :p

OpenStudy (anonymous):

the one you typed is the correct one

OpenStudy (irishboy123):

scan or link your question no point is answering to ghosts

OpenStudy (anonymous):

i don't know how to scan it and the link wont work since it's under my account name.. you already know what the problem looks like, you sent it yourself, so it should be good enough?

OpenStudy (anonymous):

\[\sqrt[4]{x^2+2x}\]

OpenStudy (anonymous):

?

OpenStudy (freckles):

what happens if you have this\[\sqrt[4]{-1}\]? does this exist over the real numbers?

OpenStudy (anonymous):

I don't think so since you shouldnt have -'s under a radical

OpenStudy (freckles):

great so you want x^2+2x to be positive or zero right?

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

for a number to be positive or equal it must be greater than or equal to 0

OpenStudy (freckles):

so you basically need to solve x^2+2x>=0

OpenStudy (freckles):

to find the domain

OpenStudy (anonymous):

do we just get rid of the radical then?

OpenStudy (freckles):

you told me just a sec ago that we didn't want the thing under the radical to be negative which means we want it to be zero or positive

OpenStudy (freckles):

the thing under the radical was x^2+2x

OpenStudy (freckles):

so you want x^2+2x>=0

OpenStudy (anonymous):

it has to be in interval notation

OpenStudy (anonymous):

so would I start with (0,_)

OpenStudy (freckles):

you would solve x^2+2x>=0

OpenStudy (freckles):

then we can put an interval notation

OpenStudy (anonymous):

x^2>or= -2x

OpenStudy (anonymous):

and then would we square both sides?? because then there'd be a negative on the right

OpenStudy (freckles):

\[x^2+2x \ge 0 \\ x(x+2) \ge 0\] draw a number line x(x+2) is zero when x=0 or x+2=0  test the intervals

OpenStudy (freckles):